# Visatronic: A Multimodal Decoder-Only Model for Speech Synthesis

### v1

@@ 2.3 Unified Multimodal Video-Text-Speech Transformer Decoder
We propose a unified multimodal decoder-only transformer architecture for processing multiple modalities – video, text and speech – in order to generate speech given video and text inputs, see Figure 1. The architecture consists of a single transformer decoder that processes the multimodal input representations from Section 2.2. Unlike traditional approaches that use one modality as input, or separate encoder(s) for multimodal input, our unified architecture enables cross-modal interactions through self-attention layers while maintaining temporal coherence. The model is trained end-to-end using cross entropy loss to predict the next discrete values in sequence, allowing it to learn intrinsic relationships across modalities that are crucial for tasks requiring multimodal understanding. During inference, the model can generate tokens autoregressively while maintaining coherence across all modalities.
@@
  • Ordering Strategy: Representations from all modalities are temporally ordered: either text, video and speech inputs; or video, text and speech inputs. For both cases, when speech is generated, the transformer decoder attends to all representations of text and video modalities. The ordering between text and video defines the interplay between them.
  • Streaming Strategy: Text tokens go first, but video and speech inputs are ordered following their original time alignment, thus preserving the natural flow of information in each of these modalities. In this approach, the speech inputs never attend to the future in time video inputs, thus reducing the sequence length processed at every speech generation step.
Positional Encoding    Due to combination of both video and text modalities for speech generation, our sequences are longer than in TTS task. Thus, capturing positional information properly is crucial. Prior work consistently showed that relative positional embeddings perform better (see e.g. [38, 3]). We apply RoPE [37] multiplicative relative positional embedding across the entire sequence. As a simple way, we maintain a global position space across all modalities, treating speaker, video, text and speech inputs uniformly in terms of positional embedding. In addition, thanks to video and speech time alignment, we investigate different position sequences to align representations appearing at the similar timestamps from different modalities, see positions notation in Figure 4.
Initialization    Placing all modalities’ inputs into one sequence for the decoder, we found that having different submodules to map each modality to the shared space leads to inconsistency of the embeddings across modalities (e.g. they have very different norm magnitudes). Thus, proper initialization of these submodules is essential. We identified that a proper scale for the initial weights distribution by bringing all inputs’ final embeddings to the same sphere is sufficient for stable and fast convergence during training.
@@
Figure 4: Input sequence layout for Visatronic. We encode all the modalities to a discrete space (see Figures 2 and 3) directly used by the decoder transformer model: each modality discrete representations are marked with its colored square. Each row represents the different strategies we adopt for combining multimodal information for learning the temporal alignment between these modalities: (top) text goes before video, while video is followed by speech; (middle) text goes first, while speech and video are ordered in time so that speech generation at time $t$ attends to the whole text but only past video at $t^{\prime}<t$; (bottom) video goes before text while text if followed by speech. Positions sequence either has global indexing across all modalities, or video and speech are aligned in time.
@@ 3 Experiments
Method Input Modality GT WER ($\downarrow$) GT (discrete) WER ($\downarrow$) WER ($\downarrow$) Sync Score ($\uparrow$) TimeSync (s) ($\downarrow$)
TTS Text 4.0 ±0.1 10.5 ±0.1 19.0$\mathord{+}8.5$
VTTS (VT-ordered) Video-Text 17.2$\mathord{+}6.7$
TTS Text 2.6 ±0.1 10.1 ±0.2 14.7$\mathord{+}4.7$ 1.54 0.62 ±0.98
[…]
VTTS (VT-ordered) Video-Text $\mathord{+}2.1$ 1.64 0.47 ±0.63
Table 1: Word error rate (WER) and TimeSync on VoxCeleb2. We report ground truth (GT) WER (computed on original audio), GT (discrete) WER (computed on the reconstructed audio from ground truth discrete speech tokens), and WER computed on generated speech. WER is calculated between ground truth text and transcription obtained by forwarding whisper-large v2 on selected audio. The first set of results (first two rows) uses PL.v1 transcriptions, while the remaining results are on PL.v2. VTTS (VT-ordered) achieves the best performance with 12.2% WER, outperforming both single-modality (TTS) and other multi-modal approaches. GT (discrete) WER (10.1±0.2%) represents the theoretical lower bound by excluding the performance loss due to speech discretization. Our goal is to minimize the gap between model WER and GT (discrete) WER, with VTTS (VT-ordered) achieving closest alignment to this lower bound with $\sim$2.1difference from its GT (discrete) WER. TimeSync shows that the video modality provides better synchronization between video and generated speech.
WER 57.4 82.4 90.6 39.2 21.4 4.5 8.2
Table 2: Generalization capability on LRS3. Word Error Rate (WER) results demonstrate strong generalization ability of our models trained on VoxCeleb2 when evaluated on LRS3 without any training on LRS3. Both VTTS (TV-ordered) and VTTS (VT-ordered) variants significantly outperform existing methods that were specifically trained on LRS3 (denoted by †).
Datasets    1) LRS3 [1] is audio-visual dataset in English, compiled from TED and TEDx video presentations. This dataset stands out for its focus on unconstrained long sentences, featuring a rich vocabulary of over 50k words and thousands of unique speakers. It contains approximately 151k videos with around 439h of speech with transcription. There are 1,452 videos in the test split. 2) VoxCeleb2 [9] is a large-scale audio-visual dataset primarily designed for speaker recognition task but applicable to various audio-visual processing domains. It consists of over 1M face-cropped YouTube videos from more than 6k distinct identities, resulting in 1.6k hours of speech w/o paired transcription. The dataset is characterized by high variability in lighting conditions, image quality, pose, and motion blur, with an average video duration of 8s. This diversity in real-world conditions makes VoxCeleb2 particularly useful for developing robust models capable of performing well in unconstrained environments. To train our models on VoxCeleb2, we first develop a pipeline for pseudo-labeling (PL) the speech using Demucs [10] for speech enhancement, Whisper-large v2 [33] for automatic transcription, and proper data filtering as the data are multilingual. The initial version of labeled data, PL.v1, was obtained by keeping English-only detected samples. Later, we improved upon it by additional filtering of inconsistent too long or too short transcriptions, leaving us with PL.v2 version of data. To evaluate our models, we randomly selected subset of 2k samples from the test set.
Objective Evaluation Metrics.    To evaluate how well generated speech preserves content information, we use the word error rate (WER) metric computed between the speech recognition model outputs from Whisper-large v2 on the audio samples and the ground truth transcripts. The synchronization score (SyncScore) is computed using the pre-trained model from [8]. This model is trained to predict the time-offset between lip crops and audio based on the distance between visual and audio embeddings over a sliding window of frames. The confidence score is computed as the difference between the median and minimum distances over this sliding window and was originally used to determine the active speaker in a multi-speaker video. From evaluation, we found that SyncScore fails in many cases and does not properly measure TTS model synchronization. For that reason, we propose a new metric, TimeSync: we take ground truth transcription and do force alignment of its phoneme sequence to each audio via HMM model from HTK [47]; the latter gives us location in time for each phoneme; finally, we compute the average absolute time difference between locations of centers of the phoneme segments for ground truth and generated audio and average across all phonemes in the test set.
Subjective Evaluation Metrics.    We randomly selected 50 samples each from VoxCeleb2 and LRS3 test data for human evaluation to assess the naturalness, intelligibility and synchronization of the generated speech following [46]. Using Mean Opinion Score (MOS) with 95% confidence intervals, human evaluators rated the speech naturalness, intelligibility and synchronization on a scale of 1 to 5, where 1 represents the worst and 5 the best quality. Details on the full protocol are provided in Appendix, Section E.
Implementation Details.   For implementation details and training configuration, refer to Appendix, Section F.
3.1 State-of-the-art Comparison
Table 1 shows a comparison of our proposed approaches and the TTS baseline trained and evaluated on VoxCeleb2 data. All results show that video brings improvement into both content generation and time synchronization.
We further evaluate on LRS3 data models which were trained on VoxCeleb2 data only. Results are shown in Table 2: VTTS (TV-ordered) achieves 4.5% WER, surpassing even LipVoicer’s 21.4% WER by a large margin, while maintaining a small gap of 2.1% from its GT (discrete) WER lower bound. These results demonstrate our models’ robust generalization to out-of-distribution data and different speaking conditions.
3.2 Human Evaluation Results
Human evaluation, presented in Tables 3 and 4, shows that VTTS (VT-ordered) achieves the best performance in Intelligibility (3.48) and Naturalness (3.20), while VTTS (TV-ordered) performs better in Synchronization (2.50). These scores approach the GT (discrete) upper bound, demonstrating the effectiveness of our proposed variants.
3.3 Ablations
Faster convergence   Table 5 shows results for different number of training steps. Our models achieve better performance at 2M iterations compared to TTS baseline, showing that video modality is speeding up training convergence.
GT 4.55 ±0.09 4.79 ±0.05 4.57 ±0.10
GT (discrete) 3.95 ±0.13 3.77 ±0.15 4.36 ±0.12
TTS 3.17 ±0.19 2.92 ±0.21 1.98 ±0.15
VTTS (TV-streaming) 3.19 ±0.17 2.99 ±0.16 2.28 ±0.17
VTTS (TV-ordered) 3.35 ±0.17 3.02 ±0.19 2.50 ±0.21
VTTS (VT-ordered) 3.48 ±0.15 3.20 ±0.19 2.48 ±0.19
Table 3: Human evaluation on VoxCeleb2. Mean Opinion Scores (MOS) (1-5) with 95% confidence intervals for Intelligibility, Naturalness, and Synchronization metrics. VTTS (VT-ordered) achieves the best performance in Intelligibility (3.48) and Naturalness (3.20), while VTTS (TV-ordered) performs best in Synchronization (2.50). Both significantly outperform TTS baseline across all metrics. Ground truth (GT) is an upper bound, while GT (discrete) is an upper bound due to speech discretization.
@@ 3.3 Ablations
Table 4: Human evaluation on LRS3. Mean Opinion Scores (MOS) (1-5) with 95% confidence intervals for our VoxCeleb2-trained models evaluated on LRS3. VTTS (TV-ordered) achieves better performance in Intelligibility (3.62) and Synchronization (3.12), while both methods perform equally well in Naturalness (3.01). Ground truth (GT) is an upper bound, while GT (discrete) is an upper bound due to speech quantization. These results demonstrate our model maintains good perceptual quality even on out-of-distribution data.
TTS 2M 2.6 ±0.1 10.1 ±0.2 17.3$\mathord{+}7.2$
VTTS (TV-ordered) 17.0$\mathord{+}6.9$
VTTS (VT-ordered) $\mathord{+}2.1$
TTS 3M 2.6 ±0.1 10.1 ±0.2 14.7$\mathord{+}4.6$
VTTS (TV-ordered) 14.1$\mathord{+}4.0$
VTTS (VT-ordered) $\mathord{+}2.1$
Table 5: Convergence analysis.   Training iterations comparison shows faster convergence when video modality is used in addition to text. While TTS requires 3M iterations, both VTTS (VT-ordered) and VTTS (TV-ordered) variants achieve comparable or better performance in only 2M iterations. VTTS (VT-ordered) demonstrates superior performance with 12.2% WER and smaller gap (+2.1%) from GT (discrete) WER, showing efficient optimization when leveraging both modalities.
@@
Different aggregation of video representations Table 6 shows results of different strategies for spatial aggregation of video representations, with simple “sum” operation performing the best.
@@
Qualitative results   Figure 5 shows mel-spectrogram comparisons between TTS, GT, and VTTS (VT-ordered). The mel-spectrogram generated by VTTS (VT-ordered) closely resembles GT in terms of temporal structure and speech patterns, particularly in capturing natural pauses and utterance duration. While TTS generates beyond the original duration (445 frames vs GT’s 393 frames) and fails to maintain proper temporal alignment, VTTS (VT-ordered) accurately matches GT’s frame length (393 frames) and successfully captures speech dynamics including pause locations. This demonstrates VTTS’s ability to leverage visual information for generating temporally coherent speech that aligns with the original video timing. The spectral patterns in VTTS (VT-ordered) also show similar energy distributions to GT, particularly in the harmonic structure during speech segments. Analysis of TimeSync for the synchronization is shown in Figures 6 and 7 for the same sample. Influence of modalities    Table 7 shows the impact of ablating individual modalities for VTTS (VT-ordered) model during evaluation. Removing the text modality severely degrades performance, leading to 74.5% WER, while removing the video modality results in 46.4% WER. These results demonstrate that both modalities contribute complementary information, highlighting the importance of our different strategies combining multimodal information.
@@
Table 6: Video inputs aggregation. Comparison of different ways to combine video frame representations before inputting into VTTS (TV-ordered) decoder. Sum operation achieves the best WER (12.2%), outperforming more complex strategies like attention (14.5%) and stacking (14.3%). This suggests that simple element-wise operations are sufficient for effective video inputs aggregation.
@@
Table 7: Effect of modality. Word Error Rate (WER) is computed when we systematically remove either text or video modality from the input for our VTTS model during evaluation only. Starting from VTTS (VT-ordered) (12.2% WER), removing text (w/o T) causes significant degradation (74.5% WER), while using only text without video (w/o V) achieves 46.4% WER. These results demonstrate the complementary nature of both modalities, with text providing more content and video significantly contributing to overall speech generation.
Figure 5: Qualitative comparison of log mel-spectrograms. Visualization of generated log mel-spectrograms: Text-to-Speech (TTS, top), Ground Truth (GT, middle), and our Video-Text-to-Speech (VTTS, VT-ordered, bottom). VTTS (VT-ordered) demonstrates better temporal alignment with GT (393 frames) compared to TTS (445 frames), showing the benefit of video conditioning for maintaining correct speech duration. The spectral patterns in VTTS (VT-ordered) also more closely match GT’s energy distribution.
Figure 7: Alignment between phonemes in ground truth (GT) speech and generated speech used for TimeSync computation. We show the time in seconds for the phoneme segment centers computed for GT (x-axis) and generated (y-axis) speech for the GT transcription. Dashed gray line is the ideal time synchronization between GT and generated speech. TTS is way out for the proper timing compared to VTTS (VT-ordered) and VTTS (TV-ordered).

### v2

@@ 2.3 Unified Multimodal Video-Text-Speech Transformer Decoder
We propose a unified multimodal decoder-only transformer architecture for processing multiple modalities – video, text and speech – in order to generate speech given video and text inputs, see Figure 1. The architecture consists of a single transformer decoder that processes discrete multimodal input representations from Section 2.2. Unlike traditional approaches that use one modality as input, or separate encoder(s) for multimodal input, our unified architecture enables cross-modal interactions through self-attention layers while maintaining temporal coherence. The model is trained end-to-end using cross entropy loss to predict the next discrete values in sequence, allowing it to learn intrinsic relationships across modalities that are crucial for tasks requiring multimodal understanding. During inference, the model can generate tokens autoregressively while maintaining coherence across all modalities.
@@
  • Ordering Strategy: All modality tokens are sorted based on temporal order – either as (text, video, speech) or (video, text, speech). In both cases, during speech generation, the decoder attends to all preceding video and text tokens, and the choice of ordering influences how cross-modal dependencies are learned.
  • Streaming Strategy: Text tokens appear at the beginning, followed by video and speech tokens ordered by their original timestamps. This preserves the temporal progression of video and speech modalities and ensures that at each decoding step, speech tokens only attend to preceding video tokens, reducing attention overhead and enforcing causal constraints.
Positional Encoding
Due to combination of both video and text modalities for speech generation, our sequences are longer than in TTS task. Thus, capturing positional information properly is crucial. Prior work consistently showed that relative positional embeddings perform better (see e.g. [47, 4]). We apply RoPE [45] multiplicative relative positional embedding across the entire sequence. As a simple way, we maintain a global position space across all modalities, treating speaker, video, text and speech inputs uniformly in terms of positional embedding. In addition, thanks to video and speech time alignment, we investigate different position sequences to align representations appearing at the similar timestamps from different modalities, see positions notation in Figure 4.
Initialization    Placing all modalities’ inputs into one sequence for the decoder, we found that having different submodules to map each modality to the shared space leads to inconsistency of the embeddings across modalities (e.g. they have very different norm magnitudes). Thus, proper initialization of these submodules is essential. We identified that a proper scale for the initial weights distribution by bringing all inputs’ final embeddings to the same sphere is sufficient for stable and fast training convergence.
@@
TimeSync
To evaluate how well the generated speech aligns temporally with the ground truth, we define a phoneme-level alignment metric, TimeSync. It is computed as:
$\texttt{TimeSync}=\frac{1}{N}\sum_{\phi^{GT}}\left|t^{model}_{f(\phi^{GT})}-t^{GT}_{\phi^{GT}}\right|$ (1)
where $N$ is the total number of phonemes $\phi^{GT}$ in the ground truth transcriptions obtained from the original audio samples. $t^{GT}_{\phi^{GT}}=(start_{\phi^{GT}}+end_{\phi^{GT}})/2$ denotes the segment’s center (in seconds) for each ground truth phoneme. $f(\phi^{GT})$ denotes the corresponding aligned phoneme in the generated audio, and $t^{model}_{f(\phi^{GT})}=(start_{f(\phi^{GT})}+end_{f(\phi^{GT})})/2$ is the center of that segment. Here, $start$ and $end$ indicate the phoneme’s start and end timestamps in either the generated or ground truth audio. This metric captures the temporal deviation between aligned phoneme centers and helps quantify how well the model preserves synchronization.
Figure 4: Input sequence for Visatronic. We encode all modalities into a discrete token space (see Figures 2 and 3), which is directly consumed by the decoder-only transformer. Each modality’s discrete representation is indicated by a colored square. Each row illustrates a different strategy for combining multimodal information to learn temporal alignment across modalities: (top) text precedes video, which is followed by speech; (middle) text appears first, while speech and video are interleaved in temporal order such that speech generation at time $t$ attends to the full text and only past video tokens at $t^{\prime}<t$; (bottom) video precedes text, which is followed by speech. The position sequence either uses global indexing across all modalities or aligns video and speech tokens by timestamp.
@@ 3 Experiments
Table 1: Human evaluation on LRS3. Mean opinion scores (MOS) (1-5) with 95% confidence intervals for our VoxCeleb2-trained models evaluated on LRS3. VTTS (TV-ordered) achieves better performance in Intelligibility (3.62) and Synchronization (3.12), while both methods perform equally well in Naturalness (3.01). GT is an upper bound, while GT (discrete) shows degradation due to speech quantization. These results demonstrate our model maintains good perceptual quality even on out-of-distribution data. Numbers or prior methods are taken directly from their respective papers.
@@
Lip to Speech Synthesis
Lip2Speech[24] 2.21 ±0.08 2.20 ±0.09 2.69 ±0.08
SVTS[36] 2.17 ±0.08 2.15 ±0.09 2.71 ±0.09
VCA-GAN[23] 2.19 ±0.08 2.20 ±0.09 2.71 ±0.08
Dubbing Methods
HPMDubbing[11] - 2.37 ±0.14 -
StyleDubber[12] - 1.79 ±0.12 -
@@
Table 2: Objective evaluation on LRS3. Word error rate (WER) results demonstrate strong generalization ability of our models trained on VoxCeleb2 when evaluated on LRS3 without any training on LRS3. Both VTTS (TV-ordered) and VTTS (VT-ordered) variants significantly outperform existing methods that were specifically trained on LRS3 (denoted by †).
WER 57.4 82.4 90.6 39.2 21.4 4.5 8.2
Datasets    1) LRS3 [1] is audio-visual dataset in English, compiled from TED and TEDx video presentations. 2) VoxCeleb2 [10] is a large-scale audio-visual dataset primarily designed for speaker recognition task but applicable to various audio-visual processing domains. Firstly, we develop a pipeline for pseudo-labeling (PL) the speech using Demucs [13] for speech enhancement, Whisper-large v2 [40] for automatic transcription, and proper data filtering as data are multilingual. The initial version of labeled data, PL.v1, was obtained by keeping English-only detected samples. Later, we improved upon it by additional filtering of inconsistent too long or too short transcriptions, resulting in a cleaner PL.v2 version. To evaluate our models, we randomly selected subset of 2k samples from the test set. Additional details are provided in Appendix.
Objective Evaluation Metrics.    To evaluate how well generated speech preserves content information, we use the word error rate (WER) metric calculated between the outputs of the Whisper-large v2 audio samples and the ground truth transcripts. To evaluate the temporal alignment between the generated and ground truth speech – and thereby indirectly assess the alignment between the video and its generated speech – we propose a new metric, TimeSync (see Section 2) computed by taking ground truth transcription and doing force alignment of its phoneme sequence to each audio via HMM model from HTK [56]; the latter gives us location in time for each phoneme; finally, we compute the average absolute time difference between locations of centers of the phoneme segments for ground truth and generated audio and average across all phonemes in the test set.
Subjective Evaluation Metrics.    We randomly selected 50 samples each from VoxCeleb2 and LRS3 test data for human evaluation to assess the naturalness, intelligibility, synchronization and emotional expressiveness of the generated speech following [55]. Using mean opinion score (MOS) with 95% confidence intervals, human evaluators rated the speech naturalness, intelligibility, synchronization and emotional expressiveness on a scale of 1 to 5, where 1 represents the worst and 5 the best quality. Details on the full protocol as well as implementation and training details are provided in Appendix.
3.1 Human Evaluation Results
Human evaluation is presented in Tables 1, where we compare our VTTS models against baselines from both lip-to-speech synthesis (Lip2Speech [24], SVTS [36], VCA-GAN [23]) and dubbing literature (HPMDubbing [11], StyleDubber [12]). Participants rated generated speech based on intelligibility, naturalness, and synchronization using 5-point mean opinion scores (MOS) with 95% confidence intervals. Our VTTS (TV-ordered) model achieves the highest scores among all: 3.62 in intelligibility and 3.12 in synchronization, and ties with VTTS (VT-ordered) on naturalness (3.01). These results are notably closer to the GT and GT (discrete) upper bounds, highlighting both the expressiveness and temporal coherence of our approach.
3.2 State-of-the-art Comparison
Table 3: WER and TimeSync on VoxCeleb2. We report three WERs: (i) GT WER from original audio, (ii) GT (discrete) WER from reconstructed audio using GT speech tokens (lower bound), and (iii) WER from model outputs. All values are computed using Whisper-large v2. The first two rows correspond to PL.v1 transcripts; the rest use PL.v2. VTTS (VT-ordered) achieves the best performance at 12.2% WER, closest to the GT (discrete) WER (10.1%) with only a $\sim$2.1gap. TimeSync further shows improved phoneme-level synchronization with video-conditioned models.
Method Input Modality GT WER ($\downarrow$) GT (discrete) WER ($\downarrow$) WER ($\downarrow$) TimeSync (s) ($\downarrow$)
TTS Text 4.0 ±0.1 10.5 ±0.1 19.0$\mathord{+}8.5$
VTTS (VT-ordered) Video-Text 17.2$\mathord{+}6.7$
TTS Text 2.6 ±0.1 10.1 ±0.2 14.7$\mathord{+}4.7$ 0.62 ±0.98
[…]
LTTS (LT-ordered) Lip-Text 14.0 $\mathord{+}3.9$ 0.46 ±0.61
VTTS (VT-ordered) Video-Text $\mathord{+}2.1$ 0.47 ±0.63
In Table 2, we further evaluate on LRS3 using models trained only on VoxCeleb2: VTTS (TV-ordered) achieves 4.5% WER, surpassing LipVoicer’s [55] 21.4% WER by a large margin, while maintaining a small gap of 2.1% from its GT (discrete) WER lower bound. These results demonstrate our model’s robust generalization to out-of-distribution data and different speaking conditions.
Table 3 shows a comparison of our proposed models and the TTS baseline trained and evaluated on VoxCeleb2. All results demonstrate that incorporating video improves both speech content generation and time synchronization. We evaluate performance using three types of WER. GT WER is computed on original audio using Whisper and serves as an ASR baseline. GT (discrete) WER is computed on audio reconstructed from ground truth discrete speech tokens using our vocoder, representing the lower bound set by quantization and vocoding artifacts. WER is the main evaluation target, measuring how well the model-generated speech aligns with the ground truth transcript. Our objective is to minimize the gap between WER and GT (discrete) WER, as this gap reflects modeling limitations beyond the quantization-vocoder pipeline. Improved annotations in PL.v2 lead to consistently better results compared to PL.v1. Furthermore, we compare our full-video VTTS model to a variant using lip crops (LTTS). VTTS achieves 12.2% WER, while LTTS gives 14.0% WER, and TTS baseline yields 14.7% WER. These results show that using full-face video provides richer information and that our model effectively leverages it without requiring lip cropping.
TimeSync.
We use https://github.com/richardbaihe/a3t from [3] to do force alignment between the phoneme sequence of ground truth transcription (PL.v2 is treated as a ground truth for VoxCeleb2) and speech: either generated or original audio. This procedure gives us phoneme location in time and phoneme duration for each audio. After, we exclude the silence (“sp”) and its duration from each alignment, see Figure 5 (left). Because every word has several possible phoneme sequences, we use Levenstein distance computation to align sequence of phonemes obtained for the generated and original audio: we consider phonemes to be aligned if they are equal or they can be obtained via substitution operation. Then, we compute the average absolute time difference between the centers of each aligned phoneme segments in generated and original audio, see Figure 5 (right).
Figure 5: TimeSync. Visualization of phoneme-level alignment used for computing the TimeSync. Left: alignment in ground truth audio before (blue) and after (green) removing silence (“sp”) segments. Right: aligned phoneme positions between ground truth (green) and generated (red) audio, where TimeSync is computed as the absolute difference between segment centers (measured in seconds).
3.3 Ablations
Different aggregation of video representations
@@ 3.3 Ablations
Table 5 shows results of different strategies for spatial aggregation of video representations before inputting into VTTS (TV-ordered) decoder, with simple “sum” operation performing the best WER (12.2%), outperforming more complex strategies like attention (14.5%) and stacking (14.3%). This suggests that simple element-wise operations are sufficient for effective video inputs aggregation.
Qualitative results
@@
Figure 6 (left) shows mel-spectrogram comparisons between TTS, GT, and VTTS (VT-ordered). The mel-spectrogram generated by VTTS (VT-ordered) closely resembles GT in terms of temporal structure and speech patterns, particularly in capturing natural pauses and utterance duration. While TTS generates beyond the original duration (445 frames vs GT’s 393 frames) and fails to maintain proper temporal alignment, VTTS (VT-ordered) accurately matches GT’s frame length (393 frames) and successfully captures speech dynamics including pause locations. This demonstrates VTTS’s ability to leverage visual information for generating temporally coherent speech that aligns with the original video timing. The spectral patterns in VTTS (VT-ordered) also show similar energy distributions to GT, particularly in the harmonic structure during speech segments. Analysis of TimeSync for the synchronization is also shown in Figure 6 (right).
Modality Drop
Table 5 shows the impact of ablating individual modalities for VTTS (VT-ordered) model during evaluation. Removing the text modality severely degrades performance, leading to 74.5% WER, while removing the video modality results in 46.4% WER. These results show that both modalities contribute complementary information, highlighting the importance of our different strategies combining multimodal inputs, with text providing more content and video significantly contributing to overall speech generation. To evaluate temporal robustness, we simulate frame drops by masking video tokens at random intervals during training/inference. This is reflected in our positional encoding strategy and token mixing logic (see Section 2.3).
Table 4: Effect of video aggregation strategies. Summation performs best.
@@
Table 5: Modality ablations. Removing text or video significantly impacts performance.
@@
Figure 6: Qualitative alignment and spectrogram analysis. Left: Log mel-spectrogram comparison for TTS (top), GT (middle), and VTTS (VT-ordered, bottom). VTTS (VT-ordered) better matches GT’s timing (393 frames) and energy patterns, unlike TTS which overextends (445 frames). Right: TimeSync visualization of phoneme alignment for the same example. Ground truth and generated phoneme segment centers are plotted on x- and y-axes respectively, with ideal sync shown as a gray diagonal. VTTS models stay closer to the diagonal than TTS, reflecting better temporal alignment.